OpenStudy (anonymous):

please help I will fan and medal A quadratic equation is shown below: 9x2 − 16x + 60 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. Part B: Solve 4x2 + 8x − 5 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used.

OpenStudy (danjs):

the radicand is the part underneath the root in the quadratic formula

OpenStudy (danjs):

b^2 - 4 a*c

OpenStudy (anonymous):

okay so how would I answer part a

OpenStudy (danjs):

the equation given has the standard form ax^2 + bx + c = 0, what is the radicand?

OpenStudy (anonymous):

would it be b

OpenStudy (danjs):

The quadratic formula is \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] the radicand under the root tells you if you have real or complex solutions depending if it is positive or negative value

OpenStudy (danjs):

b^2 - 4*a*c = (-16)^2 - 4 * 9 * 60 = -1904

OpenStudy (danjs):

a negative under a square root has imaginary solutions

OpenStudy (anonymous):

okay the as part says to Describe the solution(s) to the equation by just determining the radicand. the solutions are imaginary but how would I show my work?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

Show the calculation i did above, turns out negative

OpenStudy (danjs):

For B) not sure what appropriate method means, if it works , that is good, you may be able to factor it, but you can always apply that quadratic formula